Which of the following expressions is correct (n=no.of moles of the gas,Na=Avogadro constant,m=mass of 1 molecule of the gas,N=no.of molecules of the gas

Grade:10

6 Answers

sree lakshmi
3 years ago
Na = AVOGADRO CONSTANT =6.022×1023mol-1
m =molecular mass or mass of one molecule 
N = number of molecules 
n = number of moles 
all of them are correct 
adilakshmigogada 27
3 years ago
 
Na = AVOGADRO CONSTANT =6.022×1023mol-1
m =molecular mass or mass of one molecule 
N = number of molecules 
n = number of moles 
all of them are correct .
adilakshmigogada 27
3 years ago
 
Na = AVOGADRO CONSTANT =6.022×1023mol-1
m =molecular mass or mass of one molecule 
N = number of molecules 
n = number of moles 
all of them are correct ...
adilakshmigogada 27
3 years ago
 
Na = AVOGADRO CONSTANT =6.022×1023mol-1
m =molecular mass or mass of one molecule 
N = number of molecules 
n = number of moles 
all of them are correct ..
adilakshmigogada 27
3 years ago
 
Na = AVOGADRO CONSTANT =6.022×1023mol-1
m =molecular mass or mass of one molecule 
N = number of molecules 
n = number of moles 
all of them are correct …...
adilakshmigogada 27
3 years ago
 
Na = AVOGADRO CONSTANT =6.022×1023mol-1
m =molecular mass or mass of one molecule 
N = number of molecules 
n = number of moles 
all of them are correct